If 3x^2-2x+7=0, Then (x-1/3)^2=
Answer by checkley71(8403)

3x^2-2x+7=0
using the quadratic equation we get:
x=(2+-sqrt[-2^2+4*3*7])/2*3
x=(2+-sqrt[4-84])/6
x=(2+-sqrt-80)/6
x=(2+-8.944i)/6
x=2/6-8.944i/6
x=1/3+1.49i answer.
x=1/3-1.49i answer.
then:
(x-1/3)^2=(1/3+1.49i-1/3)^2=(1.49i)^2=-2.22 answer.
or (x-1/3)^2=(1/3-1.49i-1/3)^2=(-1.49i)^2=-2.22 answer.